OpenStudy (anonymous):

Solutions? x^2+14x+45=0? Is it x+-6 and x=2?

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

you need to factor first

OpenStudy (anonymous):

\[x^2+14x+45=(x+9)(x+4)\]

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i made a mistake anyway. it is \[(x+9)(x+5)\]

OpenStudy (anonymous):

(x+5)(x+9) so then it would be x

OpenStudy (anonymous):

x=-5 and -9

OpenStudy (anonymous):

ok got it. now you have \[(x+5)(x+9)=0\] set each factor = 0 and solve.

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

Okay cool!

OpenStudy (anonymous):

got it. good!

OpenStudy (anonymous):

but what about 4x^2+4x+1?

OpenStudy (anonymous):

now you have to factor first. this is a perfect square like the other ones you did

OpenStudy (anonymous):

So this is (2x+2)(2x+2) or 1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

2x+1 2x+1

OpenStudy (anonymous):

yes \[(2x+1)(2x+1)\]

OpenStudy (anonymous):

so this would be x=1/2?

OpenStudy (anonymous):

close

OpenStudy (anonymous):

and -1/2

OpenStudy (anonymous):

no just the second one

OpenStudy (anonymous):

both factors are the same so only one solution

OpenStudy (anonymous):

and that is the answer? -1/2?

OpenStudy (anonymous):

that is it. you have \[2x+1=0\] \[2x=-1\] \[x=-\frac{1}{2}\]

OpenStudy (anonymous):

other factor is the same so you don't need to do it again

OpenStudy (anonymous):

Okay, I think I got that down. but what if it its not perfect squares?

OpenStudy (anonymous):

then you will have two factors and perhaps 2 answers.

OpenStudy (anonymous):

perfect square gives one answer. but the problem we did before this had two answers

OpenStudy (anonymous):

so for 2x^2+7x+3 the answer would be x=-3 and -1/2?

OpenStudy (anonymous):

i don't know but i will check. did you factor?

OpenStudy (anonymous):

yes I did.

OpenStudy (anonymous):

yes you are right!

OpenStudy (anonymous):

\[(x+3)(2x+1)=0\] etc

OpenStudy (anonymous):

Great!! so for x^2-9=0 the answer will be x=3 and -3?

OpenStudy (anonymous):

yes. you can factor or just write \[x^2-9=0\] \[x^2=9\] \[x=\pm3\]

OpenStudy (anonymous):

if you factor you write \[x^2-9=(x+3)(x-3)=0\] and get the same answers

OpenStudy (anonymous):

Okay! so I just want to go over putting an equation into ax^2+bx+c=0. if it is 9-5x^2=(x+2)^2-4

OpenStudy (anonymous):

first multiply out on the right hand side

OpenStudy (anonymous):

okay done.

OpenStudy (anonymous):

then collect like terms on the right hand side. you should end up with \[x^2+4x\]

OpenStudy (anonymous):

okay. yes. I have that.

OpenStudy (anonymous):

you want this set = 0 so add \[5x^2\] to both sides and subtract 9

OpenStudy (anonymous):

you will have a 0 on the left and everything else on the rigth

OpenStudy (anonymous):

so it is 0=5x^2=4x-9? Is that the correct answer?

OpenStudy (anonymous):

no I mean. +

OpenStudy (anonymous):

between the 5x^2 and the 4x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we have \[9-5x^2=x^2+4x\]

OpenStudy (anonymous):

add \[5x^2\] to both sides to get rid of it in the left, \[9=6x^2+4x\]

OpenStudy (anonymous):

you forgot to add the \[5x^2\] to the \[x^2\]

OpenStudy (anonymous):

OH. okay. so then it is 0=6x^2+4x-9?

OpenStudy (anonymous):

then you will get \[0=6x^2+4x-9\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which of course you probably want to write as \[6x^2+4x-9=0\] not that it makes any difference

OpenStudy (anonymous):

Okay. Thanks!

OpenStudy (anonymous):

yw